# PMGS: Reconstruction of Projectile Motion across Large Spatiotemporal Spans via 3D Gaussian Splatting

### v1

@@ 3.3 Cross-modal Kalman fusion
Throughout the fusion process, we fully integrate the physical model, optical flow observations, and network outputs. The Kalman gain is used to automatically adjust the weights, thereby minimizing the error accumulation across multi-modal observation sources. Therefore, even when real-world disturbances exist or poor image quality compromises texture observations Motionblurgs ; comogs , this strategy maintains strong robustness.
4 Experiments
Datasets Source Type Phy. Mono.
LaSOT lasot Real Rigid $\times$ $\times$
D-NeRF DynaNeRF2 Syn. Def. $\times$ $\checkmark$
HyperNeRF DynaNeRF3 Real Def. $\times$ $\checkmark$
Neural3D DyNeRF Syn. Def. $\times$ $\times$
PEGASUS PEGAUS Syn. Rigid PyBullet $\checkmark$
SpringGS Spring 6S + 3R Def. Hooke’s Law $\times$
PMGS 6S + 3R Rigid Newton’s Law $\checkmark$
Grenade Missile Cola Doll Chocolate Rugby Box Shark Bear Mean
[…]
2DGS 24.82 0.895 0.153 17.42 0.714 0.325 15.97 0.847 0.195 17.99 0.792 0.218 15.19 0.758 0.275 21.56 0.830 0.185 19.72 0.973 0.172 11.57 0.794 0.101 11.78 0.823 0.201 17.34 0.825 0.202
InstantSplat 25.44 0.900 0.139 17.58 0.724 0.300 17.03 0.852 0.175 18.10 0.797 0.221 16.10 0.761 0.255 24.21 0.877 0.165 20.48 0.977 0.152 12.88 0.810 0.097 11.79 0.823 0.090 18.18 0.835 0.177
CFGS* 21.82 0.873 0.231 13.70 0.675 0.259 14.47 0.845 0.284 18.00 0.788 0.307 11.93 0.739 0.317 21.37 0.822 0.196 17.19 0.946 0.162 11.32 0.742 0.271 11.54 0.772 0.212 15.70 0.800 0.248
4DGS 20.52 0.907 0.112 22.60 0.829 0.082 20.25 0.859 0.154 23.92 0.947 0.063 17.91 0.934 0.182 Failed Failed Failed 20.19 0.874 0.131 15.32 0.709 0.302 Failed Failed Failed 19.11 0.866 0.156
MotionGS 22.64 0.876 0.166 19.14 0.857 0.156 15.39 0.823 0.217 22.70 0.854 0.119 16.99 0.902 0.115 17.37 0.855 0.151 15.92 0.512 0.464 10.38 0.654 0.353 13.99 0.887 0.231 17.59 0.782 0.227
DynamicGS 36.32 0.971 0.064 26.37 0.906 0.032 28.44 0.939 0.026 31.81 0.942 0.050 25.36 0.897 0.018 Failed Failed Failed 18.97 0.704 0.152 24.92 0.952 0.085 Failed Failed Failed 28.03 0.901 0.058
PMGS 33.58 0.967 0.011 28.87 0.915 0.025 26.50 0.940 0.014 34.38 0.959 0.014 25.01 0.924 0.020 33.72 0.954 0.012 30.45 0.974 0.019 25.07 0.952 0.094 26.64 0.959 0.059 29.35 0.949 0.029
Table 2: Comparison of video reconstruction.
Grenade Missile Cola Doll Chocolate Rugby Box Shark Bear Mean
[…]
CFGS* 0.249 0.550 0.602 0.203 0.129 0.218 0.154 0.533 0.563 0.310 0.358 0.459 0.318 0.574 0.610 0.094 0.562 0.604 0.079 0.392 0.425 0.252 0.425 0.458 0.199 0.319 0.364 0.206 0.426 0.478
4DGS 0.359 0.661 0.703 0.608 0.194 0.248 0.354 0.630 0.661 0.401 0.563 0.621 0.287 0.668 0.687 Failed Failed Failed 0.144 0.553 0.592 0.141 0.505 0.556 Failed Failed Failed 0.302 0.549 0.589
MotionGS 0.288 0.474 0.505 0.541 0.655 0.704 0.423 0.511 0.551 0.501 0.505 0.559 0.549 0.501 0.522 0.249 0.643 0.668 0.299 0.267 0.313 0.298 0.557 0.629 0.179 0.526 0.573 0.414 0.495 0.540
DynamicGS 0.548 0.227 0.269 0.910 0.118 0.156 0.898 0.182 0.226 0.663 0.311 0.363 0.995 0.070 0.090 Failed Failed Failed 0.262 0.439 0.470 0.679 0.094 0.155 Failed Failed Failed 0.707 0.205 0.247
PMGS 0.998 0.168 0.254 0.995 0.091 0.130 0.997 0.069 0.080 0.993 0.092 0.152 0.987 0.168 0.214 0.952 0.021 0.148 0.940 0.055 0.086 0.876 0.021 0.093 0.969 0.043 0.047 0.967 0.080 0.133
Table 3: Comparison of motion recovery.
4.1 Experimental Settings
Datasets. We constructed both synthetic and real-world dataset. For synthetic data, we collected 6 models from public 3D communities Blender , then set up a constant gravity field in Blender. The objects were launched with a randomly set initial velocity and accompanied by autorotation. A fixed camera was used to capture the dynamic scene at 120 FPS, rendering 120 images at a resolution of 1024x1024. For the real-world dataset, we threw 3 different objects and captured the scene using a fixed camera, with an exposure time of 1/2500s, a frame rate of 60 FPS, and a resolution of 4128×2752. During training, the images were resized to below 1600 pixels. Due to the frame rate limitations of the imaging device and the real-world projectile objects do not undergo sufficiently complete rotations (i.e., failing to exhibit a full 360° surface), we captured three projectile sequences for each model from the same viewpoint. Samples were then randomly selected from these sequences to perform modeling.
@@ 4.1 Experimental Settings
Metrics. We evaluate video reconstruction and motion recovery separately. For video reconstruction, we use PSNR, SSIM, and LPIPS SSIM ; PSNR . For motion recovery, we compute the bounding boxes of the target object in both the real and rendered image, and then calculate the IoU, absolute trajectory error (ATE) and RMSE IoU ; Tracking ; CF to comprehensively assess the spatial accuracy of tracking and trajectory. It should be clarified that, all evaluations were computed after fully removing backgrounds.
Implementation Details. We conducted experiments in the PyTorch framework with an Nvidia RTX 4090 GPU. During the centralization phase, we cropped the target from the original data and placed it at the center of a 512x512 canvas. In the registration phase, we trained a set of $[\mathbf{r,t},s]$ to align the static Gaussian with the first frame of the dynamic scene within 10,000 iterations, using a learning rate of 5e-5. This process is fast, typically taking only 2 minutes. In the motion recovery phase, the initial base learning rate applied to the frame with the minimal velocity was set to 1.0e-3, with a base iteration count of 1000. The weights for the $\mathcal{L}_{GS}$, $\mathcal{L}_{Acc}$ and $\mathcal{L}_{Smooth}$ were set to 0.7, 0.2 and 0.1.
4.2 Comparison
We select 2DGS 2DGS , InstantSplat Instantsplat as static competitors. Dynamic competitors include 4DGS 4DGS , DynamicGS Dynamic3DGS and MotionGS MotionGS . Furthermore, we employed CFGS CF as a baseline by disabling its depth estimation for reconstruction, and instead inputting our pre-trained Gaussian model for purely 6DoF pose estimation comparison (symbolized as CFGS*).
Comparison of video reconstruction. We report the quantitative results in Table 2. Our algorithm demonstrates robust performance, as the proposed density control strategy effectively enhances appearance and geometric quality, thereby providing a solid foundation for dynamic reconstruction. CFGS underperforms, primarily due to inherent flaws in depth inference pipelines. For object-specific modeling under non-open scenarios, the monocular observation system yields an underdetermined solution manifold, where maintaining 3D consistency over extended spatiotemporal intervals becomes theoretically unattainable.
Comparison of motion recovery. We uniformly sample frames from complete sequences and visualize in Figure 4, and the quantitative results are reported in Table 3. PMGS performs well in all three tracking metrics, demonstrating accurate estimation of both translation and rotation. 4DGS and MotionGS exhibit obvious trajectory fractures and artifacts—typical limitations of dynamic algorithms when modeling large-scale motions. While DynamicGS demonstrates promising performance, it exhibits limited generalization capability to certain instances. For CFGS, even with our well-trained 3D model provided, the spatial position of Gaussian is incorrectly driven toward infinity under only photometric supervision to forcibly fit image similarity. These findings collectively reveal the effectiveness of physics constraints and adaptive training scheduling for recovering complex motions across large spatiotemporal domains.
Additional experimental results and extended applications are presented in Figure 5, showing the potential of PMGS in understanding and interacting with the real world.
Figure 5: Extended experiments. (a) Novel views synthesis. (b) Multi-objects interaction. (c) Future motion trajectory prediction based on the 3D centroid coordinates with MLP. (d) Instantaneous displacement percpetion.
Video reconstruction Motion recovery
[…]
CFGS* 16.82 0.799 0.249 0.206 0.426 0.478 103 mins
Model 1 (w/o density control) 24.61 0.910 0.089 0.635 0.334 0.791 51 mins
Model 2 (w/o $L_{Acc}$ DSA) 25.63 0.929 0.045 0.886 0.137 0.188 64 mins
Model 3 (w/o Kalman) 28.95 0.940 0.035 0.933 0.098 0.147 42 mins
PMGS (Ours) 29.35 0.949 0.029 0.967 0.080 0.133 48 mins
Table 4: Ablation study of different models.
@@ 4.2 Comparison
4.3 Ablation Study
We conducted ablation experiments to verify the effectiveness of different novel modules.
@@ 4.3 Ablation Study
Effectiveness of physical constraints. Model 2 ablates $L_{Acc}$ and DSA strategy, retaining only photometric supervision and fixed training paradigm. As shown in Figure 6, the spatial position of the object exhibits drift, and errors are observed obviously in rotational orientation. The removal of DSA significantly increases the optimization time cost, revealing the necessity of adaptive adjustment of the learning rate under high-speed displacement.
Effectiveness of Kalman fusion. The removal reduced computational cost overhead at the expense of a decrease in stability. In more challenging scenarios (e.g., under substantial time-varying force interference), the strategy’s role in ensuring robustness would be more pronounced.
5 Conclusion
In this paper, we propose PMGS for reconstructing projectile motion over large spatiotemporal scales from monocular videos. Through a two-stage pipeline involving dynamic scene decomposition modeling and physically enhanced motion recovery, we effectively alleviate the issues of trajectory fragmentation and physical implausibility encountered by existing dynamic neural rendering methods when handling high-speed, non-linear rigid motion. Moreover, the proposed cross-modal Kalman fusion optimizes error accumulation from multi-source observations, further enhancing robustness against real-world disturbances. PMGS provides a novel perspective for the explicit modeling and understanding of the 3D physical world. In the future, we plan to delve into more challenging topics, such as achieving accurate 3D motion reconstruction under low-quality imaging conditions or changing force field environments.

### v3

@@ Introduction
Throughout the fusion process, we fully integrate the physical model, optical flow observations, and network outputs. The Kalman gain is used to automatically adjust the weights, thereby minimizing the error accumulation across multi-modal observation sources.
Experiments
Experimental Settings
Datasets. We constructed both synthetic and real-world datasets. For synthetic data, we collected 6 models from public 3D communities (Blender), then set up a constant gravity field in Blender. The objects were launched with a randomly set initial velocity and accompanied by autorotation. A fixed camera was used to capture the dynamic scene at 120 FPS, rendering 120 images at a resolution of 1024x1024. For the real-world dataset, we threw 3 different objects and captured the scene using a fixed camera, with an exposure time of 1/2500s, a frame rate of 60 FPS, and a resolution of 4128×2752. During training, the images were resized to below 1600 pixels. Due to the frame rate limitations of the imaging device and the real-world projectile objects do not undergo sufficiently complete rotations (i.e., failing to exhibit a full 360° surface), we captured three projectile sequences for each model from the same viewpoint. Samples were then randomly selected from these sequences to perform modeling.
@@
Metrics. We evaluate video reconstruction and motion recovery separately. For video reconstruction, we use PSNR, SSIM, and LPIPS (SSIM; PSNR). For motion recovery, we compute the bounding boxes of the target object in both the real and rendered image, and then calculate the IoU, absolute trajectory error (ATE) and RMSE (IoU; Tracking; CF) to comprehensively assess the spatial accuracy of tracking and trajectory. All evaluations are computed after fully removing backgrounds.
@@
Implementation Details. We conducted experiments in the PyTorch framework with an Nvidia RTX 4090 GPU. During the centralization phase, we cropped the target from the original data and placed it at the center of a 512x512 canvas. In the motion recovery phase, the initial base learning rate applied to the frame with the minimal velocity was set to 1.0e-3, with a base iteration count of 1000. The weights for the $\mathcal{L}_{GS}$, $\mathcal{L}_{Acc}$ and $\mathcal{L}_{Smooth}$ were set to 0.7, 0.2 and 0.1.
Comparison
We select 2DGS (2DGS), InstantSplat (Instantsplat) as static competitors. Dynamic competitors include 4DGS (4DGS), DynamicGS (Dynamic3DGS) and MotionGS (MotionGS). Furthermore, we employ CFGS (CF) as a baseline by disabling its depth estimation for reconstruction, and instead inputting our pre-trained Gaussian model for purely 6DoF pose estimation comparison (symbolized as CFGS*).
Comparison of video reconstruction. We report the quantitative results in Tables 2 and 3. Our algorithm demonstrates robust performance, as the proposed density control strategy effectively enhances appearance and geometric quality, thereby providing a solid foundation for dynamic reconstruction. CFGS underperforms, primarily due to inherent flaws in depth inference pipelines. For object-specific modeling under non-open scenarios, the monocular observation system yields an underdetermined solution manifold, where maintaining 3D consistency over extended spatiotemporal intervals becomes theoretically unattainable.
Comparison of motion recovery. We uniformly sample frames from complete sequences and visualize in Figure 4, and the quantitative results are reported in Tables 4 and 5. PMGS performs well in all three tracking metrics, demonstrating accurate estimation of both translation and rotation. 4DGS and MotionGS exhibit obvious trajectory fractures and artifacts—typical limitations of dynamic algorithms when modeling large-scale motions. While DynamicGS demonstrates promising performance, it exhibits limited generalization capability to certain instances. For CFGS, even with our well-trained 3D model provided, the spatial position is incorrectly driven toward infinity under only photometric supervision to forcibly fit image similarity. These findings collectively reveal the effectiveness of physics constraints for recovering complex motions across large spatiotemporal domains.
Ablation Study
We conduct ablation experiments to verify the effectiveness of different novel modules.
@@
Effectiveness of physical constraints. Model 2 ablates $L_{Acc}$ and the DSA strategy. As shown in Figure 5, the spatial position of the object exhibits drift, and errors are observed in rotational orientation. The removal of DSA significantly increases the optimization time cost, revealing the necessity of adaptive adjustment of the learning rate under high-speed displacement.
Effectiveness of Kalman fusion. The removal reduces computational overhead at the expense of a decrease in stability. In more challenging scenarios (e.g., under substantial time-varying force interference), the strategy’s role in ensuring robustness would be more pronounced.
Conclusion
In this paper, we propose PMGS for reconstructing projectile motion over large spatiotemporal scale from monocular videos. Through a two-stage pipeline involving dynamic scene decomposition modeling and physically enhanced motion recovery, we effectively alleviate the issues of trajectory fragmentation and physical implausibility encountered by existing dynamic neural rendering methods when handling high-speed, non-linear rigid motion. In the future, we plan to delve into more challenging topics, such as achieving accurate 3D motion reconstruction under low-quality imaging or changing force field environments.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant 82571371 and in part by the Joint Fund Project of the Ministry of Education for Equipment Pre-research under Grant 8091B042238 (Corresponding author: Chu He.), and the Fundamental Research Funds for the Central Universities Grant 204205kf0063 and the National Natural Science Foundation of China under Grant 62271354 (Corresponding author: Lei Yu.).